OpenStudy (anonymous):

I have to make a triangle that has to be 36 cm squared. What is the length?

OpenStudy (anonymous):

The length of what? I assume that 36cm squared is the area of the triangle.

OpenStudy (anonymous):

I need the total area inside. The width is 9 cm.

OpenStudy (anonymous):

What is the 36?

OpenStudy (anonymous):



OpenStudy (anonymous):

Its supposed to be the cm inside. Only, the only gave me a grid and I bottom line that is 9 cms.

OpenStudy (anonymous):

*they

OpenStudy (anonymous):

OKay so the area is 36cm squared, one of the sides is 9cm and you need to find the other side length?

OpenStudy (anonymous):

The other side is 8. Don't know if you still need it b/c closed the question. side1*side2=2*area so 9*x=2*36 9*x=72 x=8